#1.    Vista Corp. bought a piece of new equipment at the beginning of the year at...

#1.   

Vista Corp. bought a piece of new equipment at the beginning of the year at a cost of $32,500. The equipment’s estimated useful life is 4 years and the residual value is $2,500.

The estimated production that is expected from the equipment over its useful life is 100,000 units.

Expected annual production is:

Year 1             30,000 units

Year 2             40,000 units

Year 3             20,000 units

Year 4             10,000 units

a.)   Use the Units of Production depreciation method to compute the Depreciation Expense and Book Value for this new equipment for each of the four years of its useful life (your answer must include 8 amounts and the calculations to support them).

b.)   Using your answer to part a.) above, calculate the Gain or Loss that Vista Corp. would report on its income statement if it sold the equipment at the end of the third year for $ 5,300.  

Gain or Loss on sale of equipment based on Units of Production Method:
Answer #1

a)

Year Beginning book value Number of units produced Depreciation expense Ending book value
Year 1 $                           32,500 $                                    30,000 ($32,500-$2,500)/100,000*30,000 = $9,000 $                     23,500
Year 2 $                           23,500 $                                    40,000 ($32,500-$2,500)/100,000*40,000 = $12,000 $                     11,500
Year 3 $                           11,500 $                                    20,000 ($32,500-$2,500)/100,000*20,000 = $6,000 $                        5,500
Year 4 $                             5,500 $                                    10,000 ($32,500-$2,500)/100,000*10,000 = $3,000 $                        2,500
Sale value $    5,300
Less: Book value at the end of 3rd year $ (5,500)
Loss on sale of asset $     (200)

You can reach me over comment box if you have any doubts. Please rate this answer